OpenStudy (anonymous):

Which of the following is equal to the number sentence listed below? 60 + 48 A. (12 × 5)(12 × 4) B. (12 + 5)(12 + 4) C. 12 + (5 × 4) D. 12(5 + 4)

OpenStudy (anonymous):

D.

OpenStudy (pixiedust1):

Okay. 60+48=108 The only one of these is the equal to D is 12(5+4)

OpenStudy (mathstudent55):

You don't need to evaluate the expression. Start with 60 + 48, and factor 12 out of both numbers: 60 + 48 = 12 * 5 + 12 * 4 = 12(5 + 4) The answer is D.

OpenStudy (anonymous):

thanks

OpenStudy (mathstudent55):

wlcm